### v1

@@ Abstract
Sensor systems are extremely popular today and vulnerable to sensor data attacks. Due to possible devastating consequences, counteracting sensor data attacks is an extremely important topic, which has not seen sufficient study. This paper develops the first methods that accurately identify/eliminate only the problematic attacked sensor data presented to a sequence estimation/regression algorithm under a powerful attack model constructed based on known/observed attacks. The approach does not assume a known form for the statistical model of the sensor data, allowing data-driven and machine learning sequence estimation/regression algorithms to be protected. A simple protection approach for attackers not endowed with knowledge of the details of our protection approach is first developed, followed by additional processing for attacks based on protection system knowledge. In the cases tested for which it was designed, experimental results show that the simple approach achieves performance indistinguishable, to two decimal places, from that for an approach which knows which sensors are attacked. For cases where the attacker has knowledge of the protection approach, experimental results indicate the additional processing can be configured so that the worst-case degradation under the additional processing and a large number of sensors attacked can be made significantly smaller than the worst-case degradation of the simple approach, and close to an approach which knows which sensors are attacked, for the same number of attacked sensors with just a slight degradation under no attacks. Mathematical descriptions of the worst-case attacks are used to demonstrate the additional processing will provide similar advantages for cases for which we do not have numerical results. All the data-driven processing used in our approaches employ only unattacked training data.
Index Terms:
@@ I Introduction
I-A Problem and Attack Model
@@
The attack model considered in this paper allows much more powerful attacks than we have ever seen successfully detected/eliminated in the existing attack mitigation literature. We assume the attacker has complete control to generate any sensor data values after the attack. We assume the protection system has no prior knowledge of which sensors are more likely attacked. We assume the attacked sensors and the attacks launched on those sensors can change each time step. We assume the protection system can’t make any assumptions on the number of sensors attacked at any time step or about the patterns of attacks over time. Such attacks are consistent with sensor spoofing and out-of-band attacks discussed in Subsection I-A, where the signal impinging on the sensor is already modified and this modification changes a sensor output exactly as desired by the attacker. The attacked sensors can also exhibit such properties for other types of sensor attacks also, see Subsection I-A . We provide two protection approaches, one for attackers without detailed knowledge of the protection approach and another for attackers with detailed knowledge of our protection approach. We assume all the data-driven processing used in our approaches, for protection or for estimation, will employ only unattacked training data since it seems impossible to obtain training data accurately representing all possible attacks.
I-B Contribution
@@
This paper develops the first methods that accurately identify/eliminate only the problematic attacked sensor data (keep the rest) presented to a sequence estimation/regression algorithm under a proper powerful attack model that fits the situation based on known/observed attacks. Our approaches can be used as an outer shell of protection that can surround an unprotected estimation/regression algorithm (designed for unattacked sensor data), allowing the protected estimation algorithm to properly operate, with excellent performance, using the deemed to be unattacked data, which we have not seen for the attack model considered. Our approaches do not assume a known form for the statistical model of the sensor data, allowing data-driven and machine learning sequence estimation/regression algorithms to be protected. Our approaches employ only unattacked training data and handle the possible attacks allowed by the powerful attack model where the attacked sensors and attacks can change for each time step, the attacker has complete control of after-attack sensor data, and the protection system has no prior knowledge of which sensors are more likely attacked and how many sensors are attacked. Such powerful sensor attacks have not been mitigated elsewhere in the literature.
@@
We initially focus in this paper on a simple protection approach for attacks not endowed with knowledge of the details of our protection approach, but later we describe additional processing for attackers with knowledge, thus allowing a lower complexity approach if suitable. The paper provides experimental demonstration of good performance for both the simple and additional processing. Our simple method is shown to achieve, to two decimal places, indistinguishable estimation performance in cases tested, which assume the attacker does not have knowledge of the protection scheme, in comparison to an optimized approach which knows which sensors were attacked. For cases where the attacker has knowledge of the protection approach, the additional processing can be configured so that the worst-case degradation under the additional processing and a large number of sensors attacked can be made significantly smaller than the worst-case degradation under the simple approach, and close to that for an approach which knows which sensors are attacked, for the same number of attacked sensors with just a slight degradation under no attacks. Guarding against the worst-case is extremely important and if this performance can be made acceptable, then performance will always be acceptable for any attacks. Since our method does not employ high-complexity searches over all possible subsets of sensors, it is a relatively low-complexity algorithm to reduce delay and needed hardware for practical implementation.
@@ I-C Example Application
Section II provides a literature review. In Section III we describe our proposed methods and the worst-case attacks on these methods. Section IV presents our numerical results. In Section V we provide conclusions.
II Literature Review
@@
On the other hand, these EDAD approaches have an issue for the trajectory state estimation problems we consider here which typically have more than one possible (valid) state sequence. To make things clear, consider the case where the state sequence being estimated is the position/velocity of some real object. There will be many possible (valid) state sequences in such an application but many of them do not follow the position/velocity of the real object we are tracking. If the attacker substitutes sensor data following one valid sequence/trajectory with sensor data following another valid sequence/trajectory, then this attack passes EDAD. This gives the attacker tremendous power to cause big problems. The attacker can lead the system to believe the state being estimated is following a very different trajectory than it is actually following, potentially inducing an extremely large error in the estimates based on this sensor data. This error can be as large as the error between two valid trajectories that are farthest apart, unbounded if this difference is unbounded. Similar problems occur for if we estimate the state of some machine or some some other state sequence.
@@
III Proposed Methods and worst-case Attacks
@@
First, we group all sensors such that each group contains all processed sensor outputs that predict the same component of the tracked trajectory. For the group of sensors predicting the first component of the trajectory (maybe the $X$ component of an object’s position as in the top left of Fig 2), all the sensors passing EDAD (AD in Fig 2) are put through our APCC, which attempts to check if the sensor data is following the path of the actual trajectory. This same processing is carried out for each group of sensors predicting all the other components of the trajectory. Then all the sensor data that passes this check, deemed to be unattacked sensor data, will be sent on to the estimation/fusion processing which will combine (fuse) this data, possibly with other data (including trajectory estimates at previous time steps), to produce the trajectory estimate. The estimation/fusion processing can be thought of as a possibly user-selected estimation algorithm that assumes all the sensor data input to it is unattacked.
The overall approach is illustrated in the block diagram in Fig 2. It should be noted that the APCC block includes the two options just discussed. The first option, called APCC-SIMPLE, consists of just the part labeled 1 in the APCC block in Fig 2. The second option, called APCC-ADDITIONAL, includes both the parts labeled 1 and 2 (2 builds on 1) in the APCC block in Fig 2. Next, we describe the details of the APCC block, under two subsections entitled APCC-SIMPLE and APCC-ADDITIONAL. This is followed by a subsection on the worst-case attacks and a subsection providing the details of the estimation/fusion processing.
@@ III-A APCC-SIMPLE
The prediction of what a typical sensor output would look like at a given time step can employ any machine learning approach. In the numerical results, we focus on a prediction method based on the Random Forest approach as delineated by Breiman [40]. Since the number of sensors labeled as unattacked can vary over time, we employ multiple Random Forest models to accommodate different numbers of input features. Accordingly, when the number of unattacked sensors changes, a corresponding Random Forest model with the exact number of unattacked sensor inputs is selected to perform the prediction. Once the predictor is trained, the processing in APCC-SIMPLE is illustrated in Fig 3 for $\beta=99\%$.
@@
III-B APCC-ADDITIONAL
For attackers that have knowledge of our protection approach, we augment the APCC-SIMPLE approach with additional processing in an approach we call APCC-ADDITIONAL. Note that APCC-SIMPLE has already eliminated some sensors deemed to be attacked. This additional processing may eliminate more. To motivate the need for additional processing, consider the following. Suppose the malicious attackers know the details of the APCC-SIMPLE algorithm implemented to protect the sensor data. In that case, it’s reasonable to infer that the attackers might insert many attacks (called edge attacks later) just inside one edge of the confidence interval shown in Fig 3 to avoid the attacks from being identified while causing maximum damage. Thus the edge attacks are the worst-case attacks on APCC-SIMPLE. Since these attacks will not be identified without additional processing, they will skew the estimation of the trajectory. If the attacker tries to launch a large number of such attacks, our additional processing, a data-driven approach, will attempt to eliminate many (or all) of these attacks by recognizing that they are not consistent with the statistics implied by the unattacked training data. Our approach will also mitigate other attacks.
@@ III-B APCC-ADDITIONAL
In particular, APCC-ADDITIONAL uses the training data to find the smallest upper limit of a confidence interval, called $U(x)$, such that with probability $10<\alpha<$ 1 (usually closer to 1) the histogram $\hat{f}(x)$ should lie below $U(x)$ as per
@@
After learning $U(x)$ from the training data, we effectively eliminate sensor data which would cause $\hat{f}(x)$ to exceed $U(x)$. In particular, for any small histogram bin where $\hat{f}(x)>U(x)$, we exclude the smallest number of sensors that produce values in that bin so afterwards $\hat{f}(x)\leq U(x)$. The remaining sensor measurements, all those not excluded by either this or the previous processing, will be incorporated into the estimation/fusion processing as per Fig. 4 and Fig. 2.
@@
III-C Worst-case Attacks
An important point is that we can calculate both the worst-case attack performance under the simple processing in APCC-SIMPLE (the edge attack mentioned previously in Subsection III-B) and the worst-case attack performance under the additional processing in APCC-ADDITIONAL, a major contribution of our work. We will use this later in the numerical results to show the worst-case attack under APCC-ADDITIONAL can be made to cause significantly less damage that the worst-case attack under APCC-SIMPLE for cases with a large number of sensors being attacked, with the proper choice of the parameters in APCC-ADDITIONAL. If the worst-case attack can be tolerated, any attack can be tolerated.
At any given time step, the optimum $N_{A}$-sensor attack against the APCC-ADDITIONAL approach will try to insert $N_{A}$ attacks whose magnitudes lie in a set of histogram bins [^2] Bins are assumed small so we insert attacks in the middle of the bins. which each provide a large enough difference between the upper bound $U(x)$ and the sensor data histogram $\hat{f}(x)$ to support those attacks (so the attacks get through) while choosing those bins that cause the greatest degradation to the estimation/fusion. Such attacks are called water-filling attacks since one can imagine pouring water to fill up the space left between the upper bound and the sensor data histogram, but the water must be poured in the most damaging bins first.
@@ III-C Worst-case Attacks
To find the particular $N_{A}$ sensors in which we will launch those attacks, we should pick those sensors that lead to the greatest damage, called nonrandom attacks. Intuitively, these sensors will have unattacked sensor values that differ the most from the attack values. Thus if the attack will be launched on the far left-hand side of the $x$ axis as in Fig. 5 (see red bins), then you should pick a sensor whose unattacked value is the farthest away, on the right-hand side of the $x$ axis in Fig. 5 (away from the red bins).
@@
III-D Fusion Processing Block
At each time step, the possibly user-selected estimation/fusion processing will take all the sensor outputs passing the APCC and combine them with internally stored values to produce a fused scalar output for each trajectory component to be estimated. For example, the estimation/fusion processing can employ a set of previous predictions that can be saved internally inside the estimation/fusion block in Fig. 2.
The number of sensors deemed unattacked at a given time can change from time to time. This can be handled by choosing from several different machine-learning (or other) models, each with a different number of inputs. Specific examples are described in Section IV.
IV Experimental Results
First, we remind the reader that, as explained at the end of Section II, EDAD approaches alone will give very large, potentially unbounded, errors for typical sequence estimation problems. We skip numerical results of this type to save space. By definition, by adding at least APCC-SIMPL to the EDAD approaches, we can limit the errors to reasonably small values of our choosing.
As CVN data sources with a sufficient number of sensors are unavailable and to test in a fully controllable environment, we generate sensor data used for experiments from a vehicle trajectory simulated using the Simulation of Urban Mobility open source software (SUMO) and introduce controllable sensor noise, independent samples from time to time. Most experiments will be conducted for cases of both Gaussian and Laplacian sensor noise, two very different noise models [38]. Let the scale parameter of the Gaussian noise be $\sigma$ so that its variance is $\sigma^{2}$. Let the scale parameter of the Laplace noise be $b$ so that its variance is $2b^{2}$ [38]. We vary the noise variance to characterize performance for different noise levels. Here we estimate a trajectory for a scalar quantity for simplicity, where the estimation/fusion algorithm uses only the sensor data at a given time step to produce the estimate at that same time step in the trajectory (no stored data) although we have considered other cases. We first employ SUMO to generate many different possible noise-free vehicle path trajectories of length $m=150$ time steps that follow routes on a map shown in Fig 6. After adding independent noise from sensor to sensor to these trajectories, we have our sensor data. This sensor data will be used differently for testing APCC-SIMPLE and APCC-ADDITIONAL as discussed later.
@@ IV Experimental Results
None of the attacks we consider in our tests would be detected by pure EDAD since these attacks substitute one valid sequence for another as discussed in the second to last paragraph of Section II. Our purpose here is to show our new approaches can achieve very good performance for such attacks, thus providing a useful way of augmenting EDAD to perform well for attacks that EDAD misses. We should note that for the kind of attacks we consider, it’s obvious that using attacked data (rather than eliminating it) will never improve performance since these attacks delete all the useful information about the original trajectory by changing it.
@@
IV-A Numerical Results for APCC-SIMPLE: Attacks Without Knowledge
In this subsection, we discuss the numerical results obtained with the APCC-SIMPLE algorithm, described in Subsection III-A. In the experiments, the chosen $\beta$ is 99.9%, the total number of time steps in the trajectory $m$ is 150, and the sampling period of the sensor data is $1.0\times 10^{-3}$ sec. To obtain an accurate estimate of NRMSE in (3), we performed a Monte Carlo (MC) simulation where we average the NRMSE over 1 million independent realizations called the MC run length.

### v2

@@ Abstract
Due to possible devastating consequences, counteracting sensor data attacks is an extremely important topic, which has not seen sufficient study. To the best of our knowledge, this paper develops the first methods that accurately identify/eliminate only the problematic attacked sensor data presented to a sequence estimation/regression algorithm under any attack from our attack model. The approach does not assume a known form for the statistical model of the sensor data, allowing data-driven and machine learning sequence estimation/regression algorithms to be protected. A simple protection approach for attackers not endowed with knowledge of the details of our protection approach is first developed, followed by additional processing for attacks based on protection system knowledge. Experimental results show that the simple approach achieves performance indistinguishable from that for an approach which knows which sensors are attacked. For cases where the attacker has knowledge of the protection approach, experimental results indicate the additional processing can be configured so that the worst-case degradation under the additional processing and a large number of sensors attacked can be made significantly smaller than the worst-case degradation of the simple approach, and close to an approach which knows which sensors are attacked, with just a slight degradation under no attacks. Mathematical descriptions of the worst-case attacks are used to demonstrate the additional processing will provide similar advantages for cases for which we do not have numerical results. All the data-driven/machine learning processing used in our approaches employ only unattacked training data.
{IEEEkeywords}
@@ 1 Introduction
1.1 Problem and Attack Model
@@
The attack model considered in this paper allows much more powerful attacks than we have ever seen successfully detected/eliminated in the existing attack mitigation literature. We assume the attacker has complete control to generate any sensor data values after the attack. We assume the protection system has no prior knowledge of which sensors are more likely attacked. We assume the attacked sensors and the attacks launched on those sensors can change each time step. We assume the protection system can’t make any assumptions on the number of sensors attacked at any time step or about the patterns of attacks over time. Such attacks are consistent with sensor spoofing and out-of-band attacks discussed in Subsection 1.1, where the signal impinging on the sensor is already modified and this modification changes a sensor output exactly as desired by the attacker. The attacked sensors can also exhibit such properties for other types of sensor attacks also, see Subsection 1.1 . We provide two protection approaches, one for attackers without detailed knowledge of the protection approach and another for attackers with detailed knowledge of our protection approach. We assume all the data-driven/machine learning processing used in our approaches, for protection or for estimation, will employ only unattacked training data since it seems impossible to obtain training data accurately representing all possible attacks.
1.2 Contribution
@@
To the best of our knowledge, this paper develops the first methods that accurately identify/eliminate all problematic attacked sensor data (keep all the rest) presented to a sequence estimation/regression algorithm for all possible attacks in our attack model. As we discuss later in the literature review, there are other papers which make assumptions on things like the number of sensors attacked and when these assumptions are not true (yielding attacks in our attack model), these approaches provide very poor performance at the just described task. Our approaches can be used as an outer shell of protection that can surround an unprotected estimation/regression algorithm (designed for unattacked sensor data), allowing the protected estimation algorithm to properly operate, with excellent performance, using the deemed to be unattacked data for any attack in our attack model. Our approaches do not assume a known form for the statistical model of the sensor data, allowing data-driven and machine learning sequence estimation/regression algorithms to be protected. Our approaches employ only unattacked training data and handle the possible attacks allowed by our powerful attack model where the attacked sensors and attacks can change for each time step, the attacker has complete control of after-attack sensor data, and the protection system has no prior knowledge of which sensors are more likely attacked and how many sensors are attacked.
@@
We initially focus in this paper on a simple protection approach for attacks not endowed with knowledge of the details of our protection approach, but later we describe additional processing for attackers with knowledge, thus allowing a lower complexity approach if suitable. The paper provides experimental demonstration of good performance for both the simple and additional processing. Our simple method is shown to achieve nearly indistinguishable estimation performance in cases tested, which assume the attacker does not have knowledge of the protection scheme, in comparison to an optimized approach which knows which sensors were attacked. For cases where the attacker has knowledge of the protection approach, the additional processing can be configured so that the worst-case degradation under the additional processing and a large number of sensors attacked can be made significantly smaller than the worst-case degradation under the simple approach, and close to that for an approach which knows which sensors are attacked, for the same number of attacked sensors with just a slight degradation under no attacks. Guarding against the worst-case is extremely important and if this performance can be made acceptable, then performance will always be acceptable for any attacks.
@@ 1.3 Example Application
Section 2 provides a literature review. In Section 3 we describe our proposed methods and the worst-case attacks on these methods. Section 4 presents our numerical results. In Section 5 we provide conclusions.
2 Literature Review
@@
On the other hand, these EDAD approaches have an issue for the trajectory state estimation problems we consider here which typically have more than one possible (valid) state sequence. To make things clear, consider the case where the state sequence being estimated is the position/velocity of some real object. There will be many possible (valid) state sequences in such an application but many of them do not follow the position/velocity of the real object we are tracking. If the attacker substitutes sensor data following one valid sequence/trajectory with sensor data following another valid sequence/trajectory, then this attack passes EDAD. This gives the attacker tremendous power to cause big problems. The attacker can lead the system to believe the state being estimated is following a very different trajectory than it is actually following, potentially inducing an extremely large error in the estimates based on this sensor data. This error can be as large as the error between two valid trajectories that are farthest apart, unbounded if this difference is unbounded. Similar problems occur if we estimate the state of some machine or some other state sequence.
@@
3 Proposed Methods and worst-case Attacks
@@
First, we group all similar sensors such that each group contains all processed sensor outputs that predict the same component of the tracked trajectory. For the group of sensors predicting the first component of the trajectory (maybe the $X$ component of an object’s position as in the top left of Fig 2), all the sensors passing EDAD (AD in Fig 2) are put through our APCC, which attempts to check if the sensor data is following the path of the actual trajectory. This same processing is carried out for each group of sensors predicting all the other components of the trajectory. Then all the sensor data that passes this check, deemed to be unattacked sensor data, will be sent on to the estimation/fusion processing which will combine (fuse) this data, possibly with other data (including trajectory estimates at previous time steps), to produce the trajectory estimate. The estimation/fusion processing can be thought of as a possibly user-selected estimation algorithm that assumes all the sensor data input to it is unattacked.
The overall approach is illustrated in the block diagram in Fig 2. It should be noted that the APCC block includes the two options just discussed. The first option, called APCC-SIMPLE, consists of just the part labeled 1 in the APCC block in Fig 2. The second option, called APCC-ADDITIONAL, includes both the parts labeled 1 and 2 (2 builds on 1) in the APCC block in Fig 2. Next, we describe the details of the APCC block, under two subsections entitled APCC-SIMPLE and APCC-ADDITIONAL. This is followed by a subsection on the worst-case attacks.
@@ 3.1 APCC-SIMPLE
The prediction of what a typical sensor output would look like at a given time step can employ any machine learning approach. In the numerical results, we focus on a prediction method based on the Random Forest approach as delineated by Breiman [40]. Since the number of sensors labeled as unattacked can vary over time, we employ multiple Random Forest models to accommodate different numbers of input features. Accordingly, when the number of unattacked sensors changes, a corresponding Random Forest model with the exact number of unattacked sensor inputs is selected to perform the prediction. For the Random Forest models, we use a 70/10/20 train/validate/test split. Once the predictor is trained, the processing in APCC-SIMPLE is illustrated in Fig 3 for $\beta=99\%$.
@@
3.2 APCC-ADDITIONAL
For attackers that have knowledge of our protection approach, we augment the APCC-SIMPLE approach with additional processing in an approach we call APCC-ADDITIONAL. Note that APCC-SIMPLE has already eliminated some sensors deemed to be attacked. This additional processing may eliminate more. To motivate the need for additional processing, consider the following. Suppose the malicious attackers know the details of the APCC-SIMPLE algorithm implemented to protect the sensor data. In that case, it’s reasonable to infer that the attackers might insert many attacks (called edge attacks later) just inside one edge of the confidence interval shown in Fig 3 to avoid the attacks from being identified while causing maximum damage. Thus the edge attacks are the worst-case attacks on APCC-SIMPLE. Since these attacks will not be identified without additional processing, they will skew the estimation of the trajectory. If the attacker tries to launch a large number of such attacks, our additional processing will attempt to eliminate many (or all) of these attacks by recognizing that they are not consistent with the statistics implied by the unattacked training data. Our approach will also mitigate other attacks.
@@ 3.2 APCC-ADDITIONAL
In particular, APCC-ADDITIONAL uses the training data to find the smallest upper limit of a confidence interval, called $U(x)$, such that with probability $0<\alpha<1$ the histogram $\hat{f}(x)$ should lie below $U(x)$ as per
@@
After learning $U(x)$ from the training data, we effectively eliminate sensor data which would cause $\hat{f}(x)$ to exceed $U(x)$. In particular, for any histogram bin where $\hat{f}(x)>U(x)$, we exclude the smallest number of sensors that produce values in that bin so afterwards $\hat{f}(x)\leq U(x)$. The remaining sensor measurements, all those not excluded by either this or the previous processing, will be incorporated into the estimation/fusion processing as per Fig. 4 and Fig. 2.
@@
3.3 Worst-case Attacks
An important point is that we can calculate both the worst-case attack performance under the simple processing in APCC-SIMPLE (the edge attack mentioned previously in Subsection 3.2) and the worst-case attack performance under the additional processing in APCC-ADDITIONAL, a major contribution of our work. We will use this later in the numerical results to show the worst-case attack under APCC-ADDITIONAL can be made to cause significantly less damage that the worst-case attack under APCC-SIMPLE for cases with a large number of sensors being attacked, with the proper choice of the parameters in APCC-ADDITIONAL. If the worst-case attack can be tolerated, any attack can be tolerated.
At any given time step, the optimum $N_{A}$-sensor attack against the APCC-ADDITIONAL approach will try to insert $N_{A}$ attacks whose magnitudes lie in a set of histogram bins which each provide a large enough difference between the upper bound $U(x)$ and the sensor data histogram $\hat{f}(x)$ to support those attacks (so the attacks get through) while choosing those bins that cause the greatest degradation to the estimation/fusion. Such attacks are called water-filling attacks since one can imagine pouring water to fill up the space left between the upper bound and the sensor data histogram, but the water must be poured in the most damaging bins first.
@@ 3.3 Worst-case Attacks
To find the particular $N_{A}$ sensors in which we will launch those attacks, we should pick those sensors that lead to the greatest damage, called non-random attacks. Intuitively, these sensors will have unattacked sensor values that differ the most from the attack values. Thus if the attack will be launched on the far left-hand side of the $x$ axis as in Fig. 5 (see red bins), then you should pick a sensor whose unattacked value is the farthest away, on the right-hand side of the $x$ axis in Fig. 5 (away from the red bins).
@@
4 Experimental Results
First, we remind the reader that, as explained at the end of Section 2, EDAD approaches alone can yield very large, potentially unbounded, errors for typical sequence estimation problems. We provide numerical results of this type at the end of this section, where we also describe that by adding at least APCC-SIMPLE to the EDAD approaches, we can limit the errors to reasonably small values of our choosing.
As CVN data sources with a sufficient number of sensors are unavailable and to test in a fully controllable environment, we generate sensor data used for experiments from a vehicle trajectory simulated using the Simulation of Urban Mobility open source software (SUMO) and introduce controllable sensor noise, independent samples from time to time. Most experiments will be conducted for cases of both Gaussian and Laplacian sensor noise, two very different noise models [38]. Let the scale parameter of the Gaussian noise be $\sigma$ so that its variance is $\sigma^{2}$. Let the scale parameter of the Laplace noise be $b$ so that its variance is $2b^{2}$ [38]. We vary the noise variance to characterize performance for different noise levels. Here we estimate a trajectory for a scalar quantity for simplicity, where the estimation/fusion algorithm uses only the sensor data at a given time step to produce the estimate at that same time step in the trajectory (no stored data) although we have considered other cases. We first employ SUMO to generate 192 different possible noise-free vehicle path trajectories of length $m=150$ time steps that follow routes on a map shown in Fig 6. After adding independent noise from sensor to sensor to these trajectories, we have our sensor data. In the examples, the sensor data describes position. Radar is an example of this case. The noise added to each sensor has the same variance for easy interoperability. This sensor data will be used differently for testing APCC-SIMPLE and APCC-ADDITIONAL as discussed later.
@@ 4 Experimental Results
None of the attacks we consider in our tests would be detected by pure EDAD since these attacks substitute one valid sequence for another as discussed in the second to last paragraph of Section 2. Our purpose here is to show our new approaches can achieve very good performance for such attacks, thus providing a useful way of augmenting EDAD to perform well for attacks that EDAD misses. We should note that for the kind of attacks we consider, it’s obvious that using attacked data (rather than eliminating it) will never improve performance since these attacks delete all the useful information about the original trajectory by changing it.
@@
Here we describe more details on the construction procedure for $U(x)$ that we used in the numerical results. We use unattacked sensor training data, generated from SUMO with noise added. As in all machine learning processing, more training data is better to give accurate estimates of $U(x)$. In our experiments we found the approach does seem to work well with as little as 37500 labeled training data samples describing the sensor observations. This training data is processed as per Fig. 4 and Fig. 5, giving a set of data which describes many realizations of the after-histogram data. Using it, we can calculate the smallest value in each bin $x$, call it $U(x)$, such that approximately $\alpha$ percent of the total amount of data that is in the bin is below $U(x)$ for any desired $\alpha$.
We have found, through experiments, that bins of width $21\%$ as a percentage of the standard deviation of the sensor output work well. Slightly smaller bins also work well.
4.1 Numerical Results for APCC-SIMPLE: Attacks Without Knowledge
In this subsection, we discuss the numerical results obtained with the APCC-SIMPLE algorithm, described in Subsection 3.1. In the experiments, the chosen $\beta$ is 99.9%, the total number of time steps in the trajectory $m$ is 150, and the sampling period of the sensor data is $1.0\times 10^{-3}$ sec. To obtain an accurate estimate of NRMSE in (3), we performed a Monte Carlo (MC) simulation where we average the NRMSE over 1 million independent realizations called the MC run length.